# On the combinatorics of tableaux — Classification of lattices underlying Schensted correspondences

### v1

@@ Abstract.
The celebrated Robinson–Schensted algorithm and each of its variants that have attracted substantial attention can be constructed using Fomin’s “growth diagram” construction from a modular lattice that is also a weighted-differential poset. We classify all such lattices that meet certain criteria; the main criterion is that the lattice is distributive. Intuitively, these criteria seem excessively strict, but all known Fomin lattices satisfy all of these criteria, with the sole exception of one family that is not even distributive, the Young–Fibonacci lattices. Disappointingly, we discover no new Fomin lattices.
@@ 1. Introduction
The celebrated Robinson–Schensted algorithm[Rob1938a][Schen1961a][Fom1994a][Fom1995a] and each of its variants that have attracted substantial attention can be constructed[Wor2023c]*sec. 2 using Fomin’s “growth diagram” construction from a modular lattice that is also a (weighted-)differential poset[Stan1988a]. We define these requirements as a “Fomin lattice” that has a minimum element $\hat{0}$. We classify all Fomin lattices that meet certain criteria. Our main criterion is that the lattice is distributive, with secondary criteria on the set of points (the join-irreducible elements of the lattice). Intuitively, these criteria seem excessively strict, but all known Fomin lattices satisfy all of these criteria, with the sole exception of one family that is not even distributive, the Young–Fibonacci lattices. This work adds to the classes of lattices within which all of the Fomin lattices have been classified. Disappointingly, we discover no new Fomin lattices.
The history of this problem in the literature is discussed in para. 2. Para. 2 enumerates the relevant properties of structures. All known Fomin lattices are described in para. 2. They consist of Young’s lattice of partitions, the shifted Young’s lattice of strict partitions, the lattices of partitions into a limited number of parts, the Young-Fibonacci lattices, the cylindric partition lattices, the lattice of ideals of a “skew strip”, some trivial cases, and certain derivatives of other Fomin lattices. The previous classification theorems for Fomin lattices are discussed in para. 2.
Our new classification theorem is in sec. 7. The class of lattices classified by our theorem is incomparable with the classes of previous classification theorems; it does not supersede them. However, it is an incremental contribution to the classification of Fomin lattices (and thus Robinson–Schensted algorithms) and is the only one that applies to “weighted-differential” lattices, that is, where the weight on coverings is not identically 1.
Our proof techniques are largely elementary. They consist of a careful accounting of the consequences of the differential condition (7) in lattices that satisfy our criteria. This “local characterization” (sec. 5) is a succession of restatements of the differential condition (lem. 5.7, 5.13, 5.14, 5.18, 5.25, and 5.26). These results culminate in lem. 5.27, that no point covers three distinct points or is covered by three distinct points. Knowing this, the points must be placeable on a two-dimensional grid (lem. 6). We then proceed to the “global characterization” (sec. 6) of how the points can be placed on the grid. This culminates in the classification theorem 7.1.
@@ 2. Background
    Distributive lattice with further conditions (as in th. 7.1).
Significant works in the literature, and distinctive parts of them, can be classified in the following table. In some cases a work’s definitions and general analysis are considerably more general than all of the examples described in the work; in that case the classification of the definitions are marked “D” and classification of the examples are marked “E”. Some works that are largely a description of examples (often novel examples) are only marked with “E”. The works with the best previous classification results (detailed at the end of para. 2) are listed with “C” in the appropriate category.
Underlying structure Weighted Unweighted
Dual graded graph Fomin [Fom1994a]*Def. 1.3.4 Qing [Qing2008a]*sec. 2.2 E
Graded poset Stanley [Stan1988a]*sec. 2.2 DC
@@
(Modular) lattice Stanley [Stan1990a]*sec. 3
@@
    The structure is a dual graded graph (DGG): [Fom1994a] There is a graded set of vertices. In the original formulation each grade contains a finite number of vertices, but we do not include that criterion here. (But see property (D).) Vertices in a grade have edges to some vertices in adjacent ranks, forming a graph. Each edge has an assigned weight, with the weights being drawn from a set of values $V$. In the original formulation, each edge is an otherwise unweighted multi-edge, so the weights are positive integers ($V=\mathbb{Z}$), but we do not include that criterion here.
@@
    Critically, the DGG has the differential property: It has a constant $r$ called the differential degree and the DGG is called $r$-differential. Assume that weight 0 edges have been added so that both graphs have the same edges. Let $x\lessdot y$ mean that $y$ is in the rank above $x$ and $(x,y)$ is an edge. When $x\lessdot y$, let $w_{1}(x\lessdot y)$ be the weight of the edge $(x,y)$ in the first graph and $w_{2}(x\lessdot y)$ be the weight of the edge $(x,y)$ in the second graph. Then the differential property is that for every pair of vertices $x$ and $y$ in the same grade,
    $\sum_{z\,|\,x\lessdot z,y\lessdot z}w_{1}(x\lessdot z)w_{2}(x\lessdot z)-\sum_{w\,|\,w\lessdot x,w\lessdot y}w_{2}(w\lessdot x)w_{1}(w\lessdot y)=\begin{cases}0&\text{if }x\neq y\\ r&\text{if }x=y\end{cases}$ (1)
@@
    The structure is a differential poset: [Stan1988a] This is a narrower type of DGG. Both graphs have the same set of edges and the weights on the edges of the second graph are all 1. In the original formulation, the weights on the edges of the first graph are also 1, but we do not include that criterion here. Differential posets with weights that are not 1 are also called weighted-differential posets.
@@
A poset $J$ is cover-modular [^1] We choose this name because of its resemblance to the property of coverings in modular lattices. But in general, $J$ is not necessarily modular, indeed, it need not even be a lattice. A poset with this property and finite length is called a “modular poset” in [Birk1967a]*§ II.8. if when $x,y\in J$ and $x$ and $y$ both cover an element of $J$, then they are both covered by an element of $J$ (which may not be unique). Dually, we require that when $x,y\in J$ and $x$ and $y$ both are covered by an element of $J$, then they both cover an element of $J$ (which may not be unique).
@@
A poset $J$ is unique-cover-modular if when $x,y\in J$ and either $x$ and $y$ both cover an element of $J$ or $x$ and $y$ are covered by an element of $J$, then they both cover a unique element of $J$ and they are both covered by a unique element of $J$.
@@
    [Fom1994a] requires each rank of vertices to be finite. If the structure has a $\hat{0}$, this is equivalent to finite covers, but without a $\hat{0}$, it is a stronger property. (See (H).)
@@
    $\sum_{z\,|\,x\lessdot z}w(x\lessdot z)-\sum_{w\,|\,w\lessdot x}w(w\lessdot y)=r$ (2)
@@
The weights on a DGG or differential poset are projective-constant iff (for a DGG) the weight functions $w_{1}$ and $w_{2}$ or (for a differential poset) the weight function $w$ have the following property: the function $f$ is constant on coverings which are cover-projective: If $x\wedge y\lessdot x$ and $y\lessdot x\vee y$, then $f(x\wedge y\lessdot x)=f(y\lessdot x\vee y)$ and $f(x\wedge y\lessdot y)=f(x\lessdot x\vee y)$.
@@
    The structure is a differential poset that is a modular lattice: If the structure is a locally-finite modular lattice, then if the weight function is projective-constant, it is constant on projectivity classes[Birk1967a]*§V.5 of coverings.
@@
    The weighting and differential degree are admissible: However see lem. 3.2 for how we abuse the term “admissible”.
@@
A weighting $w$ and differential degree $r$ are admissible if the set of values $V$ (the target set of $w$) is $\mathbb{Z}$, all of the values of $w$ are positive, and $r$ is positive.
Depending on the degree of abstraction of the discussion, the requirements for the set of values $V$ varies. The minimum requirement is that $V$ be a module over a unitial ring of scalars SS. [^2] Remember, $V$ is a module over a ring, not a module over an algebra, so it is not assumed to have an internal product. Usually we will be unconcerned with the details of $S$ other than it necessarily contains a homomorphic image of $\mathbb{Z}$ and thus we can multiply elements of $V$ by integer coefficients. If $V=\mathbb{Z}$, then $\mathbb{Z}$ is isomorphically embedded in $S$.
@@
    The structure is a distributive lattice: In this case, we define $J$ to be the poset of points (join-irreducibles) of the lattice, so the lattice is isomorphic to the set of finite order ideals of $J$.
@@
    The poset of points $J$ is unique-cover-modular: Note that this requirement is on the poset of points, not a requirement on the structure itself (which all Fomin lattices satisfy).
@@
    The structure is a lattice that cannot be factored into two nontrivial lattices: If the lattice is distributive this is equivalent to $J$ cannot be partitioned into two disjoint non-empty posets $J=J_{1}\sqcup J_{2}$. [^3] We use $\sqcup$ for disjoint unions of sets and posets because “disjoint union” is the coproduct in the categories of sets and posets.
@@
    Any lattice satisfying (A) can be made into a degenerate Fomin lattice by setting $r=0$ and $w=0$.
@@
    [Fom1994a]*Exam. 2.1.2 and 2.2.7 Young’s lattice, $\mathbb{Y}$, the lattice of partitions, with $r=1$ and $w=1$ is an admissible Fomin lattice with $\hat{0}$.
@@
    [Sag1979b][Wor1984a][Fom1994a]*Exam. 2.2.8 The shifted Young’s Lattice, $\mathbb{SY}$, the lattice of partitions into unequal parts, with $r=1$ and $w=1\textup{ or }2$ is an admissible Fomin lattice with $\hat{0}$.
@@
    [Fom1994a]*Exam. 2.2.5 For any integer $k\geq 1$, the $k$-row Young’s lattice, $\mathbb{Y}_{k}$, the lattice of partitions into at most $k$ parts (or dually, into an arbitrary number of parts that are all $\leq k$), with $r=k$ and $w$ having a particular range of values is an admissible Fomin lattice with $\hat{0}$.
@@
    [Fom1994a]*Exam. 2.2.1 The upward semi-infinite chain $\mathbb{N}$ with $r=k$ and $w$ having a range of values (which is $\mathbb{Y}_{1}$, a special case of (4)) is an admissible Fomin lattice with $\hat{0}$.
@@
    [Stan1988a]*sec. 5[Roby1991a]*Ch. 5 For any integer $k\geq 1$, the Young–Fibonacci lattice of degree $k$, $\mathbb{YF}_{k}$, with $r=k$ and $w=1$ is an admissible Fomin lattice with $\hat{0}$.
@@
    [El2025a]*sec. 2 Fig. 2 For any integers $d,L\geq 1$, the cylindric partition lattice $\mathbb{CP}_{dL}$ with $r=0$ and $w=1$ is a Fomin lattice but is not admissible and has no $\hat{0}$.444[El2025a] uses the notation $\mathcal{L}_{d,L}$ for $\mathbb{CP}_{dL}$. The cylindric partitions have been known since at least [GessKratt1997a] but [El2025a] seems to be the first report that all of the $\mathbb{CP}_{dL}$ are 0-differential. (Though [Fom1994a] reports that the $\mathbb{CP}_{2t}$ are 0-differential.)
@@
    [Fom1994a]*Exam. 2.2.11 $\mathbb{SS}_{t}={\operatorname{Ideal}_{\textup{f}}}(t\textit{-SkewStrip})={\operatorname{Ideal}_{\textup{f}}}(\mathbb{CP}_{2t})$ with $r=0$ and $w=1\textup{ or }2$ is a Fomin lattice but is not admissible and has no $\hat{0}$. [^5] We define ${\operatorname{Ideal}_{\textup{f}}}(P)$, where $P$ is a poset, to be the lattice of finite lower order ideals of $P$.
@@
    [Fom1994a]*Lem. 2.2.3 Any cartesian product of a finite number of Fomin lattices that all have the same value set $V$, with the weight of a covering in the product lattice equal to the weight within the factor from which the covering is derived and the differential degree equal to the sum of the differential degrees of the factors, is a Fomin lattice. The product lattice has $\hat{0}$ if all factors have $\hat{0}$ and is admissible if all factors are admissible.
@@
    Given a Fomin lattice, the same lattice but multiplying the differential degree and weighting by the same constant gives a Fomin lattice. The resulting lattice is admissible if the original lattice is admissible and the constant is a positive integer.
@@
    [Byrn2012a]*Ch. 6 If a Fomin lattice with $\hat{0}$ has $r=1$ and $w=1$, then it is either $\mathbb{Y}$ or $\mathbb{YF}_{1}$.
@@
    For any $x\lessdot y$ in $L_{1}$ and $z$ in $L_{2}$, $w_{1}(x\lessdot y)=w((x,z)\lessdot(y,z))$.
@@
    For any $w$ in $L_{1}$ and $x\lessdot y$ in $L_{2}$, $w_{2}(x\lessdot y)=w((w,x)\lessdot(w,y))$.
@@
Because of property (A)(4), $w$ is constant on cover-projective pairs of elements of $L$, so equation (3) holds if $z_{1}\lessdot z_{2}$. Since any two elements $z_{1},z_{2}\in L_{2}$ can be connected by a finite chain of upward and downward covers, equation (3) holds for any $z_{1}$ and $z_{2}$. Thus the $w_{1}$ in conclusion (4) exists.
@@
We now show that $r_{1}$ and $r_{2}$ exist, that is, that $L_{1}$ is a Fomin lattice with differential degree $r_{1}$ and weighting $w_{1}$ and similarly for $L_{2}$. For any $x\in L_{1}$ and any $z_{1},z_{2}\in L_{2}$, examining $(x,z_{1})$,
@@
This shows that $\sum_{u\,|\,z\lessdot u}w_{2}(z\lessdot u)-\sum_{q\,|\,q\lessdot z}w_{2}(q\lessdot z)$ is the same for $z=z_{1}$ and $z=z_{2}$ for any $z_{1},z_{2}\in L_{2}$. We define $r_{2}$ as this value.
@@
If $L$ is a Fomin lattice with differential degree $r$ and weighting $w$, and if $L$ is the cartesian product of two Fomin lattices $L=L_{1}\times L_{2}$, with $L_{1}$ having differential degree $r_{1}$ and weighting $w_{1}$ and $L_{2}$ having differential degree $r_{2}$ and weighting $w_{2}$, then
@@
    If $L$ has a minimum element $\hat{0}_{L}$ and is admissible, and $L_{1}$ and $L_{2}$ are not the trivial lattice, then $L_{1}$ and $L_{2}$ are admissible.
@@
Regarding (2): By the construction in th. 2.7, the set of values $V_{L_{1}}$ for $L_{1}$ and the set of values $V_{L_{2}}$ for $L_{2}$ are the same as the set of values $V_{L}$ of $L$, and thus are $\mathbb{Z}$. Every value of $w_{1}$ and every value of $w_{2}$ are values of $w$ for suitable arguments, so all values of $w_{1}$ and $w_{2}$ are positive. Since $L$ has a $\hat{0}_{L}$, $L_{1}$ has a $\hat{0}_{L_{1}}$ and $L_{2}$ has a $\hat{0}_{L_{2}}$. By lem. 3.2, both $L_{1}$ and $L_{2}$ are admissible. ∎
@@
(1) A McLarnan map on $L$ may not factor into McLarnan maps on $L_{1}$ and $L_{2}$ in any nice way.
(2) The trivial lattice is a Fomin lattice: Its $w$ must be the empty function and $r=0$. It is the identity of the cartesian product of Fomin lattices. By our definition, this is not an admissible weighting. Thus lem. 2.8(2) requires the hypothesis that $L_{1}$ and $L_{2}$ are not the trivial lattice. [^7] Though our presentation of the theory might be made more uniform if we defined the trivial Fomin lattice to be admissible.
(3) Lem. 2.8(2) requires the hypothesis that $\hat{0}_{L}$ exists: Consider $\mathbb{Y}$, the lattice of partitions, which is an admissible Fomin lattice with $w=1$ and $r=1$. Consider $\mathbb{Y}^{*}$, the lattice dual of $\mathbb{Y}$, which is a Fomin lattice with $w=1$ and r=1r=-1 and so is not admissible. Define $2\mathbb{Y}$ to be the Fomin lattice $\mathbb{Y}$ with $w=2$ and $r=2$, which is admissible. Then $L=2\mathbb{Y}\times\mathbb{Y}^{*}$ is a Fomin lattice with $w=1\textup{ or }2$ and $r=1$, which is admissible. Thus $L=2\mathbb{Y}\times\mathbb{Y}^{*}$ is an admissible Fomin lattice that is a product of Fomin lattices, one of which is not admissible.

### v2

@@ Abstract.
The celebrated Robinson–Schensted algorithm and each of its variants that have attracted substantial attention can be constructed using Fomin’s “growth diagram” construction from a modular lattice that is also a weighted-differential poset. In this paper we classify all such “Fomin” lattices that meet certain criteria; the main criterion is that the lattice is distributive. Intuitively, these criteria seem excessively strict, but all known Fomin lattices satisfy all of these criteria, with the sole exception of one family that is not even distributive, the Young–Fibonacci lattices and cartesian products involving them. We discover a new class of Fomin lattices, but unfortunately they cannot be used to construct Robinson–Schensted algorithms.
@@
Note added to the 2nd version: Paragraphs 6.5, 6.6, and 6.7 of the 1st version were incorrect and have been replaced. Their conclusions remain unchanged. Also, a new class of Fomin lattices has been added as item (11) of para. 2.3: the “vertical” joining of an arbitrary doubly-infinite sequence boolean algebras $B_{1}$ or $B_{2}$.
@@ 1. Introduction
The celebrated Robinson–Schensted algorithm[Rob1938a][Schen1961a][Fom1994a][Fom1995a] and each of its variants that have attracted substantial attention can be constructed[Wor2023c]*sec. 2 using Fomin’s “growth diagram” construction from a modular lattice that is also a (weighted-)differential poset[Stan1988a]. We define these requirements as a “Fomin lattice” that has a minimum element $\hat{0}$. In this paper we classify all Fomin lattices that meet certain criteria. Our main criterion is that the lattice is distributive, with secondary criteria on the set of points (the join-irreducible elements of the lattice). Intuitively, these criteria seem excessively strict, but all known Fomin lattices satisfy all of these criteria, with the sole exception of one family that is not even distributive, the Young–Fibonacci lattices. This work adds to the classes of lattices within which all of the Fomin lattices have been classified.
We find a class of Fomin lattices that appear to be new to the literature: Let $L$ be the “vertical” joining of an arbitrary doubly-infinite sequence of 2-chains or $B_{1}$ (boolean algebras on 1 element), and diamonds or $B_{2}$ (boolean algebras on 2 elements). Then $L$ with the weighting is Fomin, but unfortunately has no $\hat{0}$. See an example in fig. 4 and para. 2(11).
The history of this problem in the literature is discussed in para. 2. Para. 2 enumerates the relevant properties of structures. All known Fomin lattices are described in para. 2. They consist of Young’s lattice of partitions, the shifted Young’s lattice of strict partitions, the lattices of partitions into a limited number of parts, the Young-Fibonacci lattices, the cylindric partition lattices, the lattice of ideals of a “skew strip”, the new class of lattices, some trivial cases, and certain derivatives of other Fomin lattices. The previous classification theorems for Fomin lattices are discussed in para. 2.
Our new classification theorem is in sec. 7. The class of lattices classified by our theorem is incomparable with the classes of previous classification theorems; it does not supersede them. However, it is an incremental contribution to the classification of Fomin lattices (and thus Robinson–Schensted algorithms) and is the only one that applies to “weighted-differential” lattices, that is, where the weight on coverings (or points) is not identically 1.
Our proof techniques are elementary. They consist of a careful accounting of the consequences of the differential condition (7) in lattices that satisfy our criteria. This “local characterization” (sec. 5) includes a succession of restatements equivalent to the differential condition (lem. 5.7, 5.17, 5.18, 5.22, 5.23, 5.27, 5.28, and 5.29). These results culminate in lem. 5.30 and 5.31, that no point covers three distinct points or is covered by three distinct points. We then proceed to the “global characterization” (sec. 6) of how the points can be placed on a two-dimensional grid. This culminates in the classification theorem, th. 7.1.
@@ 2. Background
    distributive lattice with the further conditions of th. 7.1.
Significant works in the literature, and distinctive parts of them, are classified in the following table. In some cases a work’s definitions and general analysis are considerably more general than all of the examples described in the work; in that case the classification of the definitions are marked “D” and classification of the examples are marked “E”. Some works that are largely a description of examples (often novel examples) are only marked with “E”. The works with the best previous classification results (detailed at the end of para. 2) are listed with “C” in the appropriate category.
Weighting
Weighted Unweighted
Underlying structure Dual graded graph Fomin [Fom1994a]*Def. 1.3.4 Qing [Qing2008a]*sec. 4.2 E
Graded poset Stanley [Stan1988a]*Def. 1.1 DC
@@
Lattice (modular) Stanley [Stan1990a]*sec. 3
@@
    The structure is a dual graded graph (DGG): [Fom1994a]*Def. 1.3.4 There is a graded set of vertices. In the original formulation each grade contains a finite number of vertices, but we do not include that criterion here. (But see property (D).) Vertices in a grade have edges to some vertices in adjacent grades, forming a graph. Each edge has an assigned weight, with the weights being drawn from a set of values $V$. In the original formulation, each edge is an otherwise unlabeled multi-edge, so the weights are positive integers ($V=\mathbb{Z}$), but we do not include that criterion here.
@@
    Critically, the DGG has the differential property: It has a constant $r$ called the differential degree and the DGG is called $r$-differential. Assume that weight 0 edges have been added so that both graphs have the same edges. Let $x\lessdot y$ mean that $y$ is in the grade above $x$ and $(x,y)$ is an edge. When $x\lessdot y$, let $w_{1}(x\lessdot y)$ be the weight of the edge $(x,y)$ in the first graph and $w_{2}(x\lessdot y)$ be the weight of the edge $(x,y)$ in the second graph. Then the differential property is that for every pair of vertices $x$ and $y$ in the same grade,
    $\sum_{z\,|\,x\lessdot z,y\lessdot z}w_{1}(x\lessdot z)w_{2}(x\lessdot z)-\sum_{v\,|\,v\lessdot x,v\lessdot y}w_{2}(v\lessdot x)w_{1}(v\lessdot y)=\begin{cases}0&\text{if }x\neq y\\ r&\text{if }x=y\end{cases}$ (1)
@@
    The structure is a differential poset: [Stan1988a] This is a narrower type of DGG. Both graphs have the same set of edges and the weights on the edges of the second graph are all 1. In the original formulation, the weights on the edges of the first graph are also 1, but we do not include that criterion here. (But see property (H).) Differential posets with weights that are not 1 are also called weighted-differential posets.
@@
A poset $P$ is cover-modular [^1] We choose this name because of its resemblance to the property of coverings in modular lattices. But in general, $P$ is not necessarily modular, indeed, it need not even be a lattice. A poset with this property and finite length is called a “modular poset” in [Birk1967a]*§II.8. if when $x,y\in P$ and $x$ and $y$ both cover an element of $P$, then they are both covered by an element of $P$ (which may not be unique). Dually, we require that when $x,y\in P$ and $x$ and $y$ both are covered by an element of $P$, then they both cover an element of $P$ (which may not be unique).
@@
A poset $P$ is unique-cover-modular [^2] This property is called a “uniquely modular poset” in [Proct1982a]*sec. 5. if when $x,y\in P$ and either $x$ and $y$ both cover an element of $P$ or $x$ and $y$ are covered by an element of $P$, then they both cover a unique element of $P$ and they are both covered by a unique element of $P$.
@@
    [Fom1994a]*sec. 1.1 requires each grade of vertices to be finite. If the structure has a $\hat{0}$, this is equivalent to finite covers, but without a $\hat{0}$, it is a stronger property. (See also property (H).)
@@
    $\sum_{z\,|\,x\lessdot z}w(x\lessdot z)-\sum_{v\,|\,v\lessdot x}w(v\lessdot x)=r$ (2)
@@
The weights on a DGG or differential poset are projective-constant iff (for a DGG) the weight functions $w_{1}$ and $w_{2}$ or (for a differential poset) the weight function $w$ have the following property: the function $f$ is equal on coverings which are cover-projective: If $x\vee y$ covers $x$ and $y$ and $x\wedge y$ is covered by $x$ and $y$, then $f(x\wedge y\lessdot x)=f(y\lessdot x\vee y)$ and $f(x\wedge y\lessdot y)=f(x\lessdot x\vee y)$.
@@
    The structure is a differential poset that is a modular lattice: If the structure is a locally-finite modular lattice, then if the weight function is projective-constant, it is constant on projectivity classes[Birk1967a]*§I.7 Def. of coverings.
@@
    The weighting and differential degree are positive.
@@
A weighting $w$ and differential degree $r$ (taken together) are positive if the set of values $V$ (the target set of $w$) is $\mathbb{Z}$, all of the values of $w$ are positive, and $r$ is positive.
Depending on the degree of abstraction of the discussion, the requirements for the set of values $V$ varies. The minimum requirement is that $V$ be a module over a unitial ring of scalars $S$. Usually we will be unconcerned with the details of $S$ other than it necessarily contains a homomorphic image of $\mathbb{Z}$ and thus we can multiply elements of $V$ by integer coefficients. If $V=\mathbb{Z}$, then $\mathbb{Z}$ is isomorphically embedded in $S$.
Lemma 2.7.
Proof.
In light of the above lemma, we will abuse language by using “$w$ is positive” or “the lattice is positive” to mean that the weighting and differential degree are positive, that is, $V$ is $\mathbb{Z}$, $r>0$, and all of the values of $w$ are positive.
@@
    The structure is a distributive lattice: In this case, we call the lattice $L$ and define $P$ to be the poset of points (join-irreducibles) of $L$. If the lattice is finitary, it is isomorphic to the set of finite order ideals of $P$.
@@
    The structure is a lattice and the poset of points $P$ is unique-cover-modular: Note that this requirement is on the poset of points, not a requirement on the structure itself (which all Fomin lattices satisfy).
@@
    The structure is a lattice that cannot be factored into two nontrivial lattices: If the lattice is distributive this is equivalent to that $P$ cannot be partitioned into two disjoint non-empty posets $P=P_{1}\sqcup P_{2}$. [^3] We use $\sqcup$ for disjoint unions of sets and posets because “disjoint union” is the coproduct in the categories of sets and posets.
@@
    Any lattice satisfying property (A) can be made into a degenerate Fomin lattice by setting $r=0$ and $w=0$.
@@
    [Fom1994a]*Exam. 2.1.2 and 2.2.7 Young’s lattice $\mathbb{Y}$, the lattice of partitions, with $r=1$ and $w=1$ is a positive Fomin lattice with $\hat{0}$. See fig. 1. (There are additional weightings $w$ on this lattice that make it Fomin but are not positive.)
    (a) Lattice $\mathbb{Y}$ with weights of the coverings
    (b) The quadrant $J(\mathbb{Y})$ with weights of the points
    Figure 1. Young’s lattice $\mathbb{Y}$, the lattice of partitions, with its canonical weighting.
@@
    [Sag1979b][Wor1984a][Fom1994a]*Exam. 2.2.8 The shifted Young’s Lattice $\mathbb{SY}$, the lattice of partitions into unequal parts, with $r=1$ and $w=1\textup{ or }2$ is a positive Fomin lattice with $\hat{0}$. See fig. 2.
    (a) Lattice $\mathbb{SY}$ with weights of the coverings
    (b) The octant $J(\mathbb{SY})$ with weights of the points
    Figure 2. The shifted Young’s lattice $\mathbb{SY}$, the lattice of strict partitions, with its canonical weighting.
@@
    [Fom1994a]*Exam. 2.2.5 For any integer $k\geq 1$, the $k$-row Young’s lattice, $\mathbb{Y}_{k}$, the lattice of partitions into at most $k$ parts (or dually, into an arbitrary number of parts that are all $\leq k$), with $r=k$ and $w$ having a particular range of values is a positive Fomin lattice with $\hat{0}$. See $\mathbb{Y}_{3}$ in fig. 3.
    (a) Lattice $\mathbb{Y}_{3}$ with weights of the coverings
    (b) $J(\mathbb{Y}_{3})$ with weights of the points
    Figure 3. The lattice $\mathbb{Y}_{3}$ of partitions with $\leq 3$ rows, with its canonical weighting.
@@
    [Fom1994a]*Exam. 2.2.1 The upward semi-infinite chain $\mathbb{N}$ with $r=k$ and $w(i)=k(i+1)$ (which is isomorphic to $\mathbb{Y}_{1}$, a special case of (4)) is a positive Fomin lattice with $\hat{0}$.
@@
    [Fom1994a]*Exam. 2.2.9 The infinite chain $\mathbb{Z}$ with and $w(i)=ri$ is a Fomin lattice but has no $\hat{0}$ and is not positive.
@@
    [Stan1988a]*sec. 5[Roby1991a]*Ch. 5 For any integer $k\geq 1$, the Young–Fibonacci lattice of degree $k$, $\mathbb{YF}_{k}$, with $r=k$ and $w=1$ is a positive Fomin lattice with $\hat{0}$.
  8. (8)
    [El2025a]*sec. 5 For any integers $d,L\geq 1$, the cylindric partition lattice $\mathbb{CP}_{dL}$ with $r=0$ and $w=1$ is a Fomin lattice but is not positive and has no $\hat{0}$.444[El2025a] uses the notation $\mathcal{L}_{d,L}$ for $\mathbb{CP}_{dL}$. The cylindric partitions have been known since at least [GessKratt1997a] but [El2025a] seems to be the first report that all of the $\mathbb{CP}_{dL}$ are 0-differential. See (9): [Fom1994a] reports that the $\mathbb{CP}_{2t}$ are 0-differential.
@@
    [Fom1994a]*Exam. 2.2.11 $\mathbb{SS}_{t}$ is the lattice of non-trivial ideals of $t\textit{-SkewStrip}=\mathbb{CP}_{2t}$, those that are neither empty nor all of $t\textit{-SkewStrip}$, with $r=0$ and $w=1\textup{ or }2$ is a Fomin lattice but is not positive and has no $\hat{0}$.
@@
    Let $L$ be the “vertical” joining of an arbitrary doubly-infinite sequence of 2-chains or $B_{1}$ (boolean algebras on 1 element), and diamonds or $B_{2}$ (boolean algebras on 2 elements). Then $L$ with $r=0$ and $w=1\textup{ or }2$ is a Fomin lattice but is not positive and has no $\hat{0}$. See an example in fig. 4.
    (a) A section of the poset of points, with the weights of the points
    (b) A section of the lattice, with the weights of the coverings, labeled with the generators of the ideals
    Figure 4. Example of a Fomin lattice of type (11)
@@
    [Fom1994a]*Lem. 2.2.3 Any cartesian product of a finite number of Fomin lattices that all have the same value set $V$, with the weight of a covering in the product lattice equal to the weight within the factor from which the covering is derived and the differential degree equal to the sum of the differential degrees of the factors, is a Fomin lattice. The product lattice has $\hat{0}$ if all factors have $\hat{0}$ and is positive if all factors are positive.
  13. (13)
    Given a Fomin lattice, the same lattice but multiplying the differential degree and weighting by the same constant gives a Fomin lattice. The resulting lattice is positive if the original lattice is positive and the constant is a positive integer.
@@
    [Byrn2012a]*Th. 1.3 If a Fomin lattice with $\hat{0}$ has $r=1$ and $w=1$, then it is either $\mathbb{Y}$ or $\mathbb{YF}_{1}$.
@@
    For any $x\lessdot y$ in $L_{1}$ and $z$ in $L_{2}$, $w((x,z)\lessdot(y,z))=w_{1}(x\lessdot y)$.
@@
    For any $w$ in $L_{1}$ and $x\lessdot y$ in $L_{2}$, $w((v,x)\lessdot(v,y))=w_{2}(x\lessdot y)$.
@@
Because of property (E), $w$ is constant on cover-projective pairs of elements of $L$, so equation (3) holds if $z_{1}\lessdot z_{2}$. Since any two elements $z_{1},z_{2}\in L_{2}$ can be connected by a finite chain of upward and downward covers, equation (3) holds for any $z_{1}$ and $z_{2}$. Thus the $w_{1}$ in conclusion (4) exists.
@@
We now show that $r_{1}$ and $r_{2}$ exist, that is, that $L_{1}$ is a Fomin lattice with differential degree $r_{1}$ and weighting $w_{1}$ and similarly for $L_{2}$. Consider any $x\in L_{1}$ and any $z_{1},z_{2}\in L_{2}$. Equation (2) for $(x,z_{1})$ is:
@@
This shows that $\sum_{u\,|\,z\lessdot u}w_{2}(z\lessdot u)-\sum_{q\,|\,q\lessdot z}w_{2}(q\lessdot z)$ is independent of $z\in L_{1}$. We define $r_{2}$ as this value.
@@
If $L$ is a Fomin lattice with differential degree $r$ and weighting $w$, then
@@
    If $L$ has a minimum element $\hat{0}_{L}$ and is positive, and $L_{1}$ and $L_{2}$ are not the trivial lattice, then $L_{1}$ and $L_{2}$ are positive.
@@
Regarding (2): By the construction in th. 2.8, $L_{1}$ has some differential degree $r_{1}$ and weighting $w_{1}$ and $L_{2}$ has differential degree $r_{2}$ and weighting $w_{2}$.; Every value of $w_{1}$ and every value of $w_{2}$ are values of $w$ for suitable arguments, so all values of $w_{1}$ and $w_{2}$ are elements of the set of values $V$ of $L$. Thus the set of values $V_{1}$ and $V_{2}$ of $L_{1}$ and $L_{2}$ can be considered to be $\mathbb{Z}$, and all values of $w_{1}$ and $w_{2}$ are positive. Since $L$ has a $\hat{0}_{L}$, $L_{1}$ has a $\hat{0}_{L_{1}}$ and $L_{2}$ has a $\hat{0}_{L_{2}}$. By lem. 2.7, both $r_{1}$ and $r_{2}$ are positive. Thus $L_{1}$ and $L_{2}$ are positive. ∎
@@
(1) A McLarnan map (def. 3.1(C)) on $L$ may not factor into McLarnan maps on $L_{1}$ and $L_{2}$ in any nice way.
(2) The trivial lattice is a Fomin lattice: Its $w$ must be the empty function and $r=0$. It is the identity of the cartesian product of Fomin lattices. By our definition, this is not a positive weighting. Thus lem. 2.9(2) requires the hypothesis that $L_{1}$ and $L_{2}$ are not the trivial lattice. [^6] Though our presentation of the theory might be made more uniform if we defined the trivial Fomin lattice to be positive.
(3) Lem. 2.9(2) requires the hypothesis that $\hat{0}_{L}$ exists: Consider $\mathbb{Y}$, the lattice of partitions, which is a positive Fomin lattice with $w=1$ and $r=1$. Consider $\mathbb{Y}^{*}$, the lattice dual of $\mathbb{Y}$, which is a Fomin lattice with $w=1$ and r=1r=-1 and so is not positive. Define $2\mathbb{Y}$ to be the Fomin lattice $\mathbb{Y}$ with $w=2$ and $r=2$, which is positive. Then $L=2\mathbb{Y}\times\mathbb{Y}^{*}$ is a Fomin lattice with $w=1\textup{ or }2$ and $r=1$, which is positive. Thus $L=2\mathbb{Y}\times\mathbb{Y}^{*}$ is a positive Fomin lattice that is a product of Fomin lattices, one of which is not positive.